# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
Figure S23: Average approximation ratio returned by the classical solver Gurobi, as a function of the runtime. The left-most data point corresponds to the minimum time across $1,000$ problem instances for Gurobi to return any solution at all. The right-most data point is either the runtime limit of $600$ seconds or the maximum time taken by the solver across $1,000$ problem instances to return the optimal solution. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Units: seconds.
@@
The runtime required for Gurobi to find a solution matching that of QRR* for $N\geq 256$ is orders of magnitude larger than the other solvers considered. Hence, it is computationally expensive to estimate $P$ in Eq. (S38) using multiple runs with random initial seeds, similar to what was done for simulated annealing and MQLib/Burer2002. Instead, we run Gurobi once for each of the problem instances for a maximum runtime of $600$ seconds. Over time, increasingly better solutions are found and they are tracked.
@@
In Fig. S23, we show the approximation ratio averaged over $1,000$ randomly generated problem instances for each size $N$, as a function of Gurobi’s runtime. We display the average runtime (vertical dashed line) for Gurobi to return a solution matching that of QRR* on a problem instance by problem instance basis. This runtime is equivalent to $t^{*}_{\textrm{opt}}$ in Eq. (S38). For $N\geq 2,048$, Gurobi failed to return a solution matching that of QRR* for all of the individual $1,000$ problem instances. Specifically, at $N=2,048$, Gurobi found matching solutions for only $6\%$ of the problem instances, while at $N=4,096$ it found matching solutions for $0.4\%$ of the problem instances. Therefore, we can only provide a lower bound on Gurobi’s runtime to match QRR*. We take $t^{*}_{\textrm{opt}}=600$ seconds when Gurobi fails to provide a solution matching that of QRR* within the time limit for a given problem instance. This is a lower bound that assumes that it is possible for Gurobi to yield a solution matching that of QRR* for $t^{*}_{\textrm{opt}}+\varepsilon$ with $\varepsilon\to 0$. The average $\mathbb{E}[t^{*}_{\textrm{opt}}]$ over the $1,000$ problem instances was used for Fig. 2 of the main text.
S5.10.3 Tabulated Optimal Runtime Summary
@@
In Tab. S6, we summarize the optimal runtime for three classical methods—simulated annealing, the Burer-Monteiro heuristic, and Gurobi—to match the experimental performance of the QRR* algorithm at $p=1$ on $1,000$ randomly generated problem instances. The data were extracted from Figs. S21,  S22 and  S23. For simulated annealing, we assumed knowledge of the optimal number of sweeps for estimating the optimal runtime, which would otherwise need to be searched. As such, while simulated annealing and the Burer-Monteiro algorithm display similar optimal runtimes as a function of $N$, the Burer-Monteiro solver may be considered superior.
@@
Table S6: Average optimal runtime in milliseconds (ms) for selected solvers to match the experimental performance of the QRR* algorithm at $p=1$ on $1,000$ randomly generated problem instances. The Gurobi runtime for $N\geq 2,048$ is a lower bound given that on most problem instances, the solver exceeded its time limit of $600$ seconds before finding a solution matching that of QRR*.

### v2

@@ E.9.2 Gurobi Solver
Figure 26: Average approximation ratio returned by the classical solver Gurobi, as a function of the run-time. The left-most data point corresponds to the minimum time across $1,000$ problem instances for Gurobi to return any solution at all. The right-most data point is either the run-time limit of $600$ seconds or the maximum time taken by the solver across $1,000$ problem instances to return the optimal solution. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Units: seconds.
@@
The run-time required for Gurobi to find a solution matching that of QRR* for $N\geq 256$ is orders of magnitude larger than the other solvers considered. Hence, it is computationally expensive to estimate $P$ in Eq. (41) using multiple runs with random initial seeds, similar to what was done for simulated annealing and MQLib/Burer2002. Instead, we run Gurobi once for each of the problem instances for a maximum run-time of $600$ seconds. Over time, increasingly better solutions are found and they are tracked.
@@
In Fig. 26, we show the approximation ratio averaged over $1,000$ randomly generated problem instances for each size $N$, as a function of Gurobi’s run-time. We display the average run-time (vertical dashed line) for Gurobi to return a solution matching that of QRR* on a problem instance by problem instance basis. This run-time is equivalent to $t^{*}_{\textrm{opt}}$ in Eq. (41). For $N\geq 2,048$, Gurobi failed to return a solution matching that of QRR* for all of the individual $1,000$ problem instances. Specifically, at $N=2,048$, Gurobi found matching solutions for only $6\%$ of the problem instances, while at $N=4,096$ it found matching solutions for $0.4\%$ of the problem instances. Therefore, we can only provide a lower bound on Gurobi’s run-time to match QRR*. We take $t^{*}_{\textrm{opt}}=600$ seconds when Gurobi fails to provide a solution matching that of QRR* within the time limit for a given problem instance. This is a lower bound that assumes that it is possible for Gurobi to yield a solution matching that of QRR* for $t^{*}_{\textrm{opt}}+\varepsilon$ with $\varepsilon\to 0$. The average $\mathbb{E}[t^{*}_{\textrm{opt}}]$ over the $1,000$ problem instances was used for Fig. 2 of the main text.
E.9.3 Tabulated Optimal Run-time Summary
@@
In Tab. 6, we summarize the optimal run-time for three classical methods—simulated annealing, the Burer-Monteiro heuristic, and Gurobi—to match the experimental performance of the QRR* algorithm at $p=1$ on $1,000$ randomly generated problem instances. The data were extracted from Figs. 24,  25 and  26. For simulated annealing, we assumed knowledge of the optimal number of sweeps for estimating the optimal run-time, which would otherwise need to be searched. As such, while simulated annealing and the Burer-Monteiro algorithm display similar optimal run-times as a function of $N$, the Burer-Monteiro solver may be considered superior.
@@
Table 6: Average optimal run-time in milliseconds (ms) for selected solvers to match the experimental performance of the QRR* algorithm at $p=1$ on $1,000$ randomly generated problem instances. The Gurobi run-time for $N\geq 2,048$ is a lower bound given that on most problem instances, the solver exceeded its time limit of $600$ seconds before finding a solution matching that of QRR*.